OpenStudy (anonymous):

Two balls are drawn in succession out of a box containing 5 red and 5 white balls. Find the probability that at least 1 ball was red, given that the first ball was.. A) Replaced before the second draw B) NOT replaced before the second draw a) Find the probability that at least 1 ball was red, given that the first ball was replaced before the second draw.

OpenStudy (goformit100):

I know the solution...

OpenStudy (anonymous):

Can you please help me?

OpenStudy (kropot72):

Section A) Consider the following 3 situations: (1) A: First ball red B: Second ball red The events A and B are independent therefore P(two red) = (5/10) * (5/10) ...........................(1) (2) A: First ball red B: Second ball black The events A and B are independent therefore P(one red) = (5/10) * (5/10) ...........................(2) (3) A: First ball black B: Second ball red The events A and B are independent therefore P(one red) = (5/10) * (5/10) ...........................(3) The probability that at least 1 ball was red, given that the first ball was replaced before the second draw is the sum of results (1), (2) and (3). Can you calculate?

OpenStudy (anonymous):

Sorry i can not calculate it,

OpenStudy (anonymous):

my calculator is currently not working

OpenStudy (anonymous):

so this is what it would look like then (5/10)(5/10)+(5/10)(5/10)+(5/10)(5/10)=

OpenStudy (kropot72):

You do not need a calculator to work this out: \[P(at\ least\ one\ red)=(\frac{5}{10}\times \frac{5}{10})+(\frac{5}{10}\times \frac{5}{10})+(\frac{5}{10}\times \frac{5}{10})=?\]

OpenStudy (kropot72):

Hint: \[\frac{5}{10}\times \frac{5}{10}=\frac{1}{2}\times \frac{1}{2}\]

OpenStudy (anonymous):

so the answer would be 0.75

OpenStudy (kropot72):

Good work! Your answer is correct.

OpenStudy (anonymous):

oops no sorry my mistake

OpenStudy (kropot72):

0.75 or 3/4 is correct.

OpenStudy (anonymous):

How would i find the probability that at least 1 ball was red, given that the first ball s not replaced before the second draw?

OpenStudy (anonymous):

take the chance after by subtracting out already drawn balls

OpenStudy (anonymous):

hmm?

OpenStudy (kropot72):

Section B) Probability first ball is red = 5/10 Probability second ball is red = 4/9 P(two red) = (5/10 * 4/9) ................(1) Probability first ball is red = 5/10 Probability second ball is black = 5/9 P(one red) = (5/10 * 5/9) ...............(2) Probability first ball is black = 5/10 Probability second ball is red = 5/9 P(one red) = (5/10 * 5/9) ...............(3) \[P(at\ least\ one\ red)=(\frac{1}{2}\times \frac{4}{9})+(\frac{1}{2}\times \frac{5}{9})+(\frac{1}{2}\times \frac{5}{9})=?\]

OpenStudy (anonymous):

thank you and the answer i have recieved was 0.77777777777

OpenStudy (anonymous):

is that also what you got?

OpenStudy (kropot72):

Your answer to B) is correct. Good work :)

OpenStudy (anonymous):

thank you for your help! :)

OpenStudy (anonymous):

but should i put my answer as 0.78?

OpenStudy (anonymous):

it says to simply my answer into a integer or fraction

OpenStudy (anonymous):

simplify**

OpenStudy (anonymous):

oops nvm its 7/9 hahaha thanks!

OpenStudy (kropot72):

To put the answer as a fraction just add the following fractions and simplify: \[\frac{4}{18}+\frac{5}{18}+\frac{5}{18}=?\]

OpenStudy (kropot72):

Yes, 7/9 is correct!

OpenStudy (anonymous):

:)

OpenStudy (kropot72):

You're welcome :)